OpenStudy (anonymous):

The pH of a chemical solution is given by the formula pH = -log10[H+] where [H+] is the concentration of hydrogen ions in moles per liter. Find the pH if the [H+] = 7.7 × 10-2.

OpenStudy (anonymous):

= log ( 7.7.10^(-2) )

OpenStudy (anonymous):

think there is a minus sign out front

OpenStudy (anonymous):

Oh negative !

OpenStudy (anonymous):

yes for some reason

OpenStudy (anonymous):

thank you guys so much :)

OpenStudy (anonymous):

\[-\log(7.7\times 10^{-2})=-\log(7.7)+2\]

OpenStudy (anonymous):

as \(-\log(7.7\times 10^{-2})=-(\log(7.7)-\log(10^{-2}))=-\log(7.7)+2\log(10)\) \(=\log(7.7)+2\)